There are 250 dogs at a dog show that weigh an average of 12 pounds,with a standard deviation of 8 pounds.if 4 dogs are chosen at random what is the probability that the average weight is greater than 8 pounds?

1
Expert's answer
2022-03-22T04:19:00-0400

"\mu=12, \sigma=8, n=4"

"P(X>8)=P(z>\frac{8-12}{\frac{8}{\sqrt{n}}})=P(Z>-1)"

"=1-P(Z<-1)=1-0.1587=0.8413"